# Time-IMM: A Dataset and Benchmark for Irregular Multimodal Multivariate Time Series

### v2

@@ Abstract
Time series data in real-world applications such as healthcare, climate modeling, and finance are often irregular, multimodal, and messy, with varying sampling rates, asynchronous modalities, and pervasive missingness. However, existing benchmarks typically assume clean, regularly sampled, unimodal data, creating a significant gap between research and real-world deployment. We introduce Time-IMM, a dataset specifically designed to capture cause-driven irregularity in multimodal multivariate time series. Time-IMM represents nine distinct types of time series irregularity, categorized into trigger-based, constraint-based, and artifact-based mechanisms. Complementing the dataset, we introduce IMM-TSF, a benchmark library for forecasting on irregular multimodal time series, enabling asynchronous integration and realistic evaluation. IMM-TSF includes specialized fusion modules, including a timestamp-to-text fusion module and a multimodality fusion module, which support both recency-aware averaging and attention-based integration strategies. Empirical results demonstrate that explicitly modeling multimodality on irregular time series data leads to substantial gains in forecasting performance. Time-IMM and IMM-TSF provide a foundation for advancing time series analysis under real-world conditions. The dataset is publicly available at https://www.kaggle.com/datasets/blacksnail789521/time-imm/data, and the benchmark library can be accessed at https://anonymous.4open.science/r/IMMTSF_NeurIPS2025.
@@ 2.1 Taxonomy of Irregularity in Time Series
Figure 2: Visualization of three representative features from each dataset in Time-IMM, illustrating the diverse sampling patterns and distinct types of irregularity.
This taxonomy highlights that real-world irregularities are not caused by a single factor but arise from fundamentally different sources, each requiring tailored modeling strategies to capture their structure and implications. Each dataset in Time-IMM exemplifies a distinct type of irregularity, as shown in Figure 2. To complement the taxonomy, Figure 2 presents representative visualizations of three features per dataset, illustrating the diversity of sampling patterns and irregular behaviors. Table 1 summarizes dataset statistics in Time-IMM. We also introduce a set of metrics to characterize time series irregularity—feature observability entropy, temporal observability entropy, and mean inter-observation interval—which respectively quantify: (1) the variability of missingness across features via the normalized Shannon entropy of the feature-presence distribution, (2) the dispersion of events over time using the normalized Shannon entropy of the event-time distribution, and (3) the average time gap between successive observations. Formal definitions of these metrics are provided in Appendix C.
@@ 2.1.1 Trigger-Based Irregularities
Trigger-based irregularities are characterized by observations that occur only in response to discrete events or system triggers. We further categorize this group into three representative types, each illustrating different real-world mechanisms leading to event-driven sampling patterns.
Event-Based Logging captures scenarios where observations occur only when notable external events arise, such as geopolitical protests or seismic activity. To represent this pattern, we collect event-driven time series from the GDELT Global Database gdelt_paper , where each entry corresponds to a discrete global event like a political crisis or major policy change.
Adaptive or Reactive Sampling describes cases where the system dynamically adjusts its sampling rate based on observed activity levels, such as environmental monitoring systems increasing sampling during severe weather or networks logging more frequently during high congestion periods. We illustrate this in our benchmark with RepoHealth, a dataset collected from GitHub project histories where sampling intervals adaptively shorten during periods of high commit activity and lengthen during quieter phases.
Human-Initiated Observations involve data recording triggered by human judgment or manual action, such as doctors measuring patient vitals based on clinical necessity or users reporting symptoms only when they perceive significant changes. To capture this behavior, we include samples from the MIMIC clinical database mimicIV ; MimicNote , where vital signs like heart rate and blood pressure are recorded irregularly according to clinicians’ assessments of patient needs.
@@ 2.1.2 Constraint-Based Irregularities
Constraint-based irregularities occur when limitations on resources, schedules, or human factors restrict when data can be collected, shaping the sampling pattern according to external constraints rather than an intended schedule. We further categorize this group into three representative types that reflect how operational and resource factors introduce structured irregularities.
Operational Window Sampling refers to situations where observations are possible only during predefined operational periods, such as stock exchanges being open only during business hours or satellites collecting imagery only during orbital passes. To represent this type of irregularity, we construct datasets from FNSPIDfnspid , where financial time series are naturally sampled only during market operating hours, leading to regular gaps outside trading periods.
Resource-Aware Collection describes cases where data is sampled conditionally based on system activity, often to reduce overhead or avoid redundant logging. In large-scale compute clusters, for example, telemetry may be collected only while jobs are running, leading to activity-dependent irregularity. We illustrate this with the ClusterTrace datasetclustertrace , which captures GPU cluster activity at Alibaba over eight weeks. Here, machine-level metrics are recorded only when containers are active, producing structured gaps tied to workload presence and resource scheduling.
Human Scheduling / Availability involves data collection governed by human activity schedules, such as smartphone sensor logs being recorded primarily when users are awake, or clinical measurements being taken during scheduled patient rounds rather than continuously. To illustrate this phenomenon, we include data from the StudentLifestudentlife study, where smartphone sensing data is inherently irregular due to human behavioral patterns and daily schedules.
@@ 2.1.3 Artifact-Based Irregularities
Missing Data / Gaps occur when expected data points are not recorded, leaving empty time steps in the sequence. These gaps often result from incomplete submissions, system errors, or data quality issues. In public health settings, missing values can arise when clinics do not submit reports or when technical problems affect data collection. We use the ILINet dataset time_mmd from the CDC to demonstrate this type of irregularity. While ILINet provides weekly records of influenza-like illness across the U.S., some weeks contain missing or partial data due to non-reporting providers or processing issues.
@@ 2.2 Dataset Construction
To ensure the textual modality is both relevant and informative, we apply a preprocessing strategy inspired by Time-MMD time_mmd . This includes an initial filtering step to retain only semantically relevant documents, followed by summarization to improve clarity and compress dense information into concise, interpretable content. Summarization is performed using GPT-4.1 Nano openai_gpt41_nano with prompt instructions tailored to preserve factual grounding, avoid hallucination, and allow abstention when relevance is uncertain. Prompt templates and preprocessing examples are included in Appendix E and Appendix F.
@@
Unlike traditional multimodal datasets time_mmd ; multimodalmimic that assume or enforce temporal alignment between modalities, Time-IMM explicitly preserves asynchronous timestamping. Numerical time series and textual observations are allowed to occur on independent, irregular schedules. We do not impose alignment between the timestamps of numerical and textual data—this reflects real-world conditions where modalities are generated through fundamentally different processes (e.g., continuous sensors vs. episodic reports). This design choice enables more flexible modeling, supporting realistic multimodal fusion strategies that can account for both recency and content semantics.
@@ 3.1 Problem Formulation
To incorporate unstructured textual information, we extend the irregular TSF setup to a multimodal setting. In addition to $X$, we observe a sequence of time-stamped textual data $S=\{(\tau_{j},s_{j})\}_{j=1}^{L_{S}}$, where $s_{j}$ is a text input observed at time $\tau_{j}\in[T_{\text{start}},T_{\text{cut}}]$, and $L_{S}$ denotes the number of text observations. Here, $T_{\text{start}}$ is the shared start time for both the time series and the text modality, and no alignment is assumed between the time series timestamps $\{t_{i}^{(n)}\}$ and the text timestamps $\{\tau_{j}\}$. The multimodal forecasting objective is to learn a function $g_{\theta}:(X,S,Q)\mapsto Y$ that integrates the temporal dynamics of the irregular time series and contextual signals from the text to produce accurate predictions $Y=\{[\hat{y}_{k}^{(n)}]_{k=1}^{F_{n}}\}_{n=1}^{N}$ at the query times $Q\subset(T_{\text{cut}},T_{\text{end}}]$.
@@ 3.2 Multimodal TSF Library
IMM-TSF supports both irregular time series models and regular forecasting architectures adapted to irregular settings. To adapt regular time series models for irregular data, we follow the canonical pre-alignment strategy introduced in t-PatchGNN t_patchgnn . This method interpolates irregular data to a fixed temporal grid, treating relative time and missingness masks as additional input variables. Since forecast queries remain part of the input in our problem formulation, we append their timestamps to the pre-aligned sequence during inference, enabling the model to condition predictions on the desired query times. Further implementation details are provided in Appendix H.
@@ 4.2 Results and Analysis
Forecasting difficulty in the unimodal setting varies substantially across different subtypes of time series irregularity. Figure 6(a) plots per-dataset forecasting errors across baselines, grouped by the irregularity subtypes defined in our taxonomy. Datasets characterized by Missing Data (ILINet), Resource-Aware Collection (ClusterTrace), or Event-Based Logging (GDELT) are notably more difficult, as their observation patterns are sparse, erratic, or driven by latent semantic triggers that are inherently harder to model. In contrast, datasets such as RepoHealth (Adaptive Sampling), FNSPID (Operational Window Sampling), and EPA-Air (Multi-Source Asynchrony) tend to be easier to model due to more predictable or recoverable observation structures. For instance, in the FNSPID dataset, which involves stock price forecasting, the irregularity stems from known business-hour trading windows—an expected pattern that simplifies temporal reasoning.
@@
To evaluate the design choices within our fusion framework, we conduct ablations over the Timestamp-to-Text Fusion (TTF) and Multimodality Fusion (MMF) modules, as summarized in Figure 6(b). For TTF, we observe no notable performance difference between the Time2Vec-augmented cross-attention (T2V-XAttn) and the simpler recency-based averaging (RecAvg), suggesting that both strategies for handling irregular timestamps in textual data are comparably effective. For MMF, GRU-gated residual addition (GR-Add) consistently outperforms cross-attention addition (XAttn-Add), suggesting that learnable gating provides a more effective means of integrating text-derived context into time series representations. By selectively controlling the flow of information, the gating mechanism in GR-Add helps suppress noisy or less relevant textual signals, leading to more stable and robust fusion.
@@
We evaluate the impact of different LLM backbones on forecasting performance in the irregular multimodal TSF setting by varying the text encoder within IMM-TSF. Specifically, we test GPT-2, BERT, LLaMA 3.1 (8B), and DeepSeek (7B), and assess their performance across representative datasets. As shown in Figure 6(c), we observe no significant performance differences across these models, suggesting that the choice of LLM backbone does not strongly influence results under irregular conditions. This lack of sensitivity may reflect the nature of the irregular TSF task: because forecasting hinges more on temporal alignment and contextual anchoring than on deep semantic reasoning, increasing LLM capacity offers no clear benefit. Additionally, while our TTF and MMF modules enable integration of textual signals into the forecasting pipeline, the overall scale of our datasets remains relatively small compared to the large-scale data used to pretrained LLMs, which may limit the advantages of larger models in this setting.
@@ 5 Future Work
Our current benchmark focuses on unstructured text as the secondary modality, but many real-world applications involve other asynchronous inputs such as images, audio, or tabular data. Extending irregular time series datasets to support multimodal integration beyond text remains an important direction for building truly general purpose time series models.
@@ Appendix A Related Dataset Work
Recent progress in time series analysis has been supported by a growing number of publicly available benchmarks ts_bench_1 ; ts_bench_2 ; ts_bench_3 ; ts_bench_4 . However, most existing datasets make simplified assumptions that limit their realism and utility for modeling complex real-world phenomena. In particular, prior time series datasets often fall short along several key dimensions that Time-IMM is explicitly designed to address.
Lack of Irregularity. Many standard time series datasets assume regularly sampled data where all features are observed at fixed intervals informer ; uci_electricity_data ; caltrans_traffic_data ; ncei_weather_data . Even in multimodal contexts, existing benchmarks often maintain regular sampling, smoothing over the irregular behaviors common in real-world systems time_mmd ; ts_survey . This assumption neglects the complexities found in domains like healthcare, environmental sensing, and finance, where sampling can be triggered by events, human decisions, or operational constraints mugrep ; health_monitoring ; gru_ode .
Lack of Asynchronous Textual Modality. Most existing time series datasets include only numerical observations and lack any unstructured textual modality ucr_archive ; m4 . While some benchmarks incorporate alternative modalities such as images, audio, or video lrs3 ; wildfirets ; multimodal_survey_1 ; multimodal_survey_2 , this work focuses specifically on unstructured textual data as the complementary modality to time series. Yet in many domains, textual signals—such as clinical notes, system logs, or policy descriptions—are crucial for interpretability and context-aware forecasting log_forecasting ; clinic_note_predicting ; policy_forecasting . Even in datasets that include both time series and text time_mmd ; multimodalmimic , the modalities are often artificially aligned by enforcing shared timestamps, disregarding the fact that different modalities in real-world settings typically follow inherently asynchronous temporal patterns. Our work explicitly supports asynchronous timestamps across modalities, preserving their natural temporal structure for more realistic modeling.
Lack of Reasoning About Irregularity Causes. While some prior datasets contain irregularly sampled time series t_patchgnn ; physionet ; activity ; ushcn , they do not provide any structured explanation or annotation of why the irregularity occurs. Irregular sampling is typically treated as a nuisance to be smoothed or interpolated away, rather than a meaningful signal that could inform model design or interpretation ts_imputation_1 ; ts_imputation_2 ; ts_imputation_3 . This lack of causal reasoning limits model robustness and interpretability, especially in high-stakes domains such as healthcare, environmental monitoring, or fault diagnostics environment_monitoring ; coke . In contrast, our work introduces a structured taxonomy of irregularity, categorizing it into trigger-based, constraint-based, and artifact-based causes to support more interpretable and realistic modeling.

### v3

@@ Abstract
Time series data in real-world applications such as healthcare, climate modeling, and finance are often irregular, multimodal, and messy, with varying sampling rates, asynchronous modalities, and pervasive missingness. However, existing benchmarks typically assume clean, regularly sampled, unimodal data, creating a significant gap between research and real-world deployment. We introduce Time-IMM, a dataset specifically designed to capture cause-driven irregularity in multimodal multivariate time series. Time-IMM represents nine distinct types of time series irregularity, categorized into trigger-based, constraint-based, and artifact-based mechanisms. Complementing the dataset, we introduce IMM-TSF, a benchmark library for forecasting on irregular multimodal time series, enabling asynchronous integration and realistic evaluation. IMM-TSF includes specialized fusion modules, including a timestamp-to-text fusion module and a multimodality fusion module, which support both recency-aware averaging and attention-based integration strategies. Empirical results demonstrate that explicitly modeling multimodality on irregular time series data leads to substantial gains in forecasting performance. Time-IMM and IMM-TSF provide a foundation for advancing time series analysis under real-world conditions. The dataset is publicly available at https://www.kaggle.com/datasets/blacksnail789521/time-imm/data, and the benchmark library can be accessed at https://github.com/blacksnail789521/IMM-TSF.
@@ 2.1 Taxonomy of Irregularity in Time Series
Figure 2: Visualization of the Time-IMM datasets, annotated to show sampling patterns linked to the taxonomy, such as adaptive bursts (RepoHealth), trading-hour gaps (FNSPID), and sensor misalignment (EPA-Air).
This taxonomy highlights that real-world irregularities are not caused by a single factor but arise from fundamentally different sources, each requiring tailored modeling strategies to capture their structure and implications. Each dataset in Time-IMM exemplifies a distinct type of irregularity, as shown in Figure 2. To complement the taxonomy, Figure 2 presents representative visualizations of three features per dataset, illustrating the diversity of sampling patterns and irregular behaviors. Table 1 summarizes dataset statistics in Time-IMM. We also introduce a set of metrics to characterize time series irregularity—feature observability entropy, temporal observability entropy, and mean inter-observation interval—which respectively quantify: (1) the variability of missingness across features via the normalized Shannon entropy of the feature‐presence distribution, (2) the dispersion of events over time using the normalized Shannon entropy of the event‐time distribution, and (3) the average time gap between successive observations. Formal definitions of these metrics are provided in Appendix C.
@@ 2.1.1 Trigger-Based Irregularities
Trigger-based irregularities are characterized by observations that occur only in response to discrete events or system triggers. We further categorize this group into three representative types, each illustrating distinct mechanisms behind event-driven sampling.
Event-Based Logging occurs when observations are recorded only upon notable external events, such as geopolitical protests or seismic activity. We illustrate this pattern using the GDELT Global Database gdelt_paper , where each record corresponds to a discrete global event like a political crisis or major policy change.
Adaptive or Reactive Sampling arises when a system automatically adjusts its sampling rate based on observed dynamics, such as a heart monitor increasing frequency during arrhythmia or a network logger reacting to high congestion. Here, the triggering is system-driven. In our benchmark, the RepoHealth dataset reflects this behavior, with GitHub project histories sampled more densely during periods of high commit activity and sparsely during quiet phases.
Human-Initiated Observations occur when data are recorded by deliberate human action rather than automation—for instance, clinicians measuring vitals only when clinically necessary or users self-reporting symptoms upon noticeable changes. This type is human-driven, distinguishing it from the automatic reactivity of adaptive sampling. We represent it using the MIMIC clinical database mimicIV ; MimicNote , where vital signs are logged irregularly according to clinicians’ judgment.
@@ 2.1.2 Constraint-Based Irregularities
Constraint-based irregularities occur when external limitations such as operational schedules, resource restrictions, or human factors determine when data can be collected. While all share the characteristic of externally imposed constraints, they differ in their predictability and therefore require distinct modeling strategies. We categorize them into three subtypes reflecting this spectrum from highly predictable to human-driven variability.
Operational Window Sampling is the most predictable form, where data collection is limited to fixed time windows (for example, trading hours or orbital passes). These patterns follow rigid schedules and can often be modeled using periodic masks or known calendars. We illustrate this type using FNSPID fnspid , where financial time series are sampled only during market operating hours, creating regular, clock-based gaps.
Resource-Aware Collection represents a partly predictable case, where sampling adapts to system resource conditions such as power, bandwidth, or workload. The resulting gaps are conditionally triggered but can often be anticipated using auxiliary signals (for example, device status). This is exemplified by ClusterTrace clustertrace , where GPU cluster telemetry is recorded only when containers are active, yielding structured but dynamic irregularity.
Human Scheduling / Availability is the least predictable form, driven by human activity patterns such as staff shifts, sleep cycles, or daily routines. These soft, behavior-dependent gaps benefit from time-aware modeling using cues like time of day or day of week. We illustrate this with StudentLife studentlife , where smartphone sensing data exhibits irregularity shaped by human schedules and activity cycles.
@@ 2.1.3 Artifact-Based Irregularities
Unplanned Missing Data / Gaps occur when expected data points are unintentionally lost due to accidents, system failures, or transmission errors, leaving empty time steps in the sequence. Unlike missingness in other categories such as adaptive sampling, human-initiated logging, or multi-source asynchrony, which is intentional or structural, these gaps arise without design or decision. They are typically treated as random missingness (MCAR or MAR) and handled with imputation, whereas intentional gaps are often informative (MNAR) and can carry modeling value. We illustrate this form of irregularity using the ILINet dataset time_mmd from the CDC, which provides weekly influenza-like illness reports across the United States. Some weeks contain missing or partial data caused by non-reporting providers or technical failures during data transmission.
@@ 2.2 Dataset Construction
To ensure the textual modality is both relevant and informative, we apply a preprocessing strategy inspired by Time-MMD time_mmd . This includes an initial filtering step to retain only semantically relevant documents, followed by summarization to improve clarity and compress dense information into concise, interpretable content. Summarization and filtering are performed jointly using GPT-4.1 Nano openai_gpt41_nano with domain-specific prompts that generate a five-sentence summary only if the input is topically relevant, otherwise skipping the document entirely. Details on how semantically relevant documents are identified, along with the full prompt templates and filtering logic, are provided in Appendix E, and representative preprocessing examples are shown in Appendix F. We use summarized textual data instead of original documents because many source datasets are proprietary and cannot be redistributed under the NeurIPS CC BY 4.0 requirement, following the same practice as Time-MMD time_mmd .
@@
Unlike traditional multimodal datasets time_mmd ; multimodalmimic that assume or enforce temporal alignment between modalities, Time-IMM intentionally preserves asynchronous timestamping. Numerical and textual observations occur on independent, irregular schedules, reflecting real-world conditions where continuous sensors and episodic reports are generated through different processes. Rather than forcing alignment, which can be ambiguous or misleading under irregular sampling, our design allows models to learn soft temporal associations through the timestamp-to-text fusion module described in Section 3.2. This approach retains full flexibility for downstream work to apply alternative alignment strategies, such as window-based or delay-aware matching, without constraining the data itself.
@@ 3.1 Problem Formulation
To incorporate unstructured textual information, we extend the irregular TSF setup to a multimodal setting. In addition to $X$, we observe a sequence of time-stamped textual data $S=\{(\tau_{j},s_{j})\}_{j=1}^{L_{S}}$, where $s_{j}$ is a text input observed at time $\tau_{j}\in[T_{\text{start}},T_{\text{cut}}]$, and $L_{S}$ denotes the number of text observations. Here, $T_{\text{start}}$ is the shared start time for both the time series and the text modality, and no alignment is assumed between the time series timestamps $\{t_{i}^{(n)}\}$ and the text timestamps $\{\tau_{j}\}$. The multimodal forecasting objective is to learn a function $g_{\theta}:(X,S,Q)\mapsto Y$ that integrates the temporal dynamics of the irregular time series and contextual signals from the text to produce accurate predictions $Y=\{[\hat{y}_{k}^{(n)}]_{k=1}^{F_{n}}\}_{n=1}^{N}$ at the query times $Q\subset(T_{\text{cut}},T_{\text{end}}]$. An overview of this multimodal irregular forecasting setup is illustrated in Figure 4.
@@ 3.2 Multimodal TSF Library
IMM-TSF supports both irregular time series models and regular forecasting architectures adapted to irregular data. To adapt regular models without imputing away irregularity, we apply a canonical pre-alignment strategy following t-PatchGNN t_patchgnn . Pre-alignment converts variable-length, unevenly spaced inputs into fixed-length tensors while preserving timestamps and missingness masks, allowing models to learn directly from irregular observation patterns. Specifically, observed values, binary masks, and normalized timestamps are aligned to a global temporal grid, and future query times are appended as unobserved points to guide inference. This approach offers a transparent, assumption-free way to reuse existing forecasting architectures under irregular sampling, unlike Gaussian Process–based methods that perform full imputation before modeling. Further algorithmic details and an illustrative example are provided in Appendix H.
@@ 4.1 Experimental Setup
To further examine temporal robustness, we conduct rolling-origin cross-validation to evaluate generalization across different time windows. Models trained on Time-IMM show stable forecasting performance under temporal drift, demonstrating strong generalization across time. Detailed results are provided in Appendix P.
@@ 4.2 Results and Analysis
Forecasting difficulty in the unimodal setting varies substantially across different subtypes of time series irregularity. Figure 6(a) plots per-dataset forecasting errors across baselines, grouped by the irregularity subtypes defined in our taxonomy. Datasets characterized by Unplanned Missing Data (ILINet), Resource-Aware Collection (ClusterTrace), or Event-Based Logging (GDELT) are notably more difficult, as their observation patterns are erratic or driven by latent semantic triggers that are inherently harder to model. In contrast, datasets such as RepoHealth (Adaptive Sampling), FNSPID (Operational Window Sampling), and EPA-Air (Multi-Source Asynchrony) tend to be easier to model due to more predictable or recoverable observation structures. For instance, in the FNSPID dataset, which involves stock price forecasting, the irregularity stems from known business-hour trading windows—an expected pattern that simplifies temporal reasoning.
@@
To evaluate the design choices within our fusion framework, we conduct ablations over the Timestamp-to-Text Fusion (TTF) and Multimodality Fusion (MMF) modules, as summarized in Figure 6(b). For TTF, we observe no notable performance difference between the Time2Vec-augmented cross-attention (T2V-XAttn) and the simpler recency-based averaging (RecAvg), suggesting that both strategies for handling irregular timestamps in textual data are comparably effective. For MMF, GRU-gated residual addition (GR-Add) consistently outperforms cross-attention addition (XAttn-Add), suggesting that learnable gating provides a more effective means of integrating text-derived context into time series representations. By selectively controlling the flow of information, the gating mechanism in GR-Add helps suppress noisy or less relevant textual signals, leading to more stable and robust fusion. A detailed comparison of computational cost and efficiency across different TTF and MMF module configurations is provided in Appendix N.1.
@@
We evaluate the impact of different LLM backbones on forecasting performance in the irregular multimodal TSF setting by varying the text encoder within IMM-TSF. Specifically, we test GPT-2, BERT, Llama 3.1 (8B), and DeepSeek (7B), and assess their performance across representative datasets. As shown in Figure 6(c), we observe no significant performance differences across these models, suggesting that the choice of LLM backbone does not strongly influence results under irregular conditions. This lack of sensitivity may reflect the nature of the irregular TSF task: because forecasting hinges more on temporal alignment and contextual anchoring than on deep semantic reasoning, increasing LLM capacity offers no clear benefit. Additionally, while our TTF and MMF modules enable integration of textual signals into the forecasting pipeline, the overall scale of our datasets remains relatively small compared to the large-scale data used to pretrain LLMs, which may limit the advantages of larger models in this setting. Further experiments, including a lightweight encoder trained from scratch (Doc2Vec) and evaluations with very large LLMs (for example, Llama-70B), are presented in Appendix M, while a detailed analysis of text encoder efficiency is provided in Appendix N.2.
@@ 5 Future Work
Our current benchmark focuses on unstructured text as the secondary modality, but many real-world applications involve other asynchronous inputs such as images, audio, or tabular data. Extending irregular time series datasets to support multimodal integration beyond text remains an important direction for building truly general-purpose time series models. As a proof of concept, we demonstrate in Appendix O that IMM-TSF can be readily extended to incorporate image-based modalities without architectural changes, highlighting its flexibility as a foundation for future multimodal extensions.
Compositional Irregularities.
Each dataset in Time-IMM currently represents a single, well-defined irregularity pattern to ensure interpretability and controlled benchmarking. A natural next step is to support datasets that combine multiple irregularity sources, reflecting the complex conditions found in real-world systems. We are developing APIs for compositional irregularity generation, allowing users to specify and mix atomic patterns (for example, human scheduling with missing data) to create hybrid datasets with richer temporal dynamics and more challenging forecasting scenarios.
@@ Appendix A Related Dataset Work
Recent progress in time series analysis has been supported by a growing number of publicly available benchmarks [67, 43, 52, 59]. However, most existing datasets make simplified assumptions that limit their realism and utility for modeling complex real-world phenomena. In particular, prior time series datasets often fall short along several key dimensions that Time-IMM is explicitly designed to address.
Lack of Irregularity. Many standard time series datasets assume regularly sampled data where all features are observed at fixed intervals [81, 55, 49, 21]. Even in multimodal contexts, existing benchmarks often maintain regular sampling, smoothing over the irregular behaviors common in real-world systems [41, 38]. This assumption neglects the complexities found in domains like healthcare, environmental sensing, and finance, where sampling can be triggered by events, human decisions, or operational constraints [77, 73, 15].
Lack of Asynchronous Textual Modality. Most existing time series datasets include only numerical observations and lack any unstructured textual modality [14, 44]. While some benchmarks incorporate alternative modalities such as images, audio, or video [1, 3, 25, 71], this work focuses specifically on unstructured textual data as the complementary modality to time series. Yet in many domains, textual signals—such as clinical notes, system logs, or policy descriptions—are crucial for interpretability and context-aware forecasting [8, 17, 7]. Even in datasets that include both time series and text [41, 79], the modalities are often artificially aligned by enforcing shared timestamps, disregarding the fact that different modalities in real-world settings typically follow inherently asynchronous temporal patterns. Our work explicitly supports asynchronous timestamps across modalities, preserving their natural temporal structure for more realistic modeling.
Lack of Reasoning About Irregularity Causes. While some prior datasets contain irregularly sampled time series [78, 60, 62, 45], they do not provide any structured explanation or annotation of why the irregularity occurs. Irregular sampling is typically treated as a nuisance to be smoothed or interpolated away, rather than a meaningful signal that could inform model design or interpretation [30, 64, 74]. This lack of causal reasoning limits model robustness and interpretability, especially in high-stakes domains such as healthcare, environmental monitoring, or fault diagnostics [4, 51]. In contrast, our work introduces a structured taxonomy of irregularity, categorizing it into trigger-based, constraint-based, and artifact-based causes to support more interpretable and realistic modeling.